### v1

# Continuous-wave, high-resolution, ultra-broadband mid-infrared nonlinear spectroscopy with tunable plasmonic nanocavities

Abstract
Vibrational sum- and difference-frequency generation (SFG and DFG) spectroscopy is a technique to probe the nonlinear response of interfaces at mid-infrared wavelengths while detecting the upconverted signal at visible wavelengths. It has recently transitioned from large-area molecular films and colloidal suspensions to nanoscale structures, enabled by dual-resonant plasmonic nanocavities that confine light and matter in the same deep-subwavelength sensing volume. Here, we implement continuous-wave ultra-broadband SFG and DFG spectroscopy from to on a variety of dual-resonant antennas under ambient conditions. The use of a commercial, broadly tunable quantum-cascade laser and the absence of geometric phase-matching requirements speed up and simplify data acquisition. At the same time, simultaneous access to SFG and DFG sidebands permits ratiometric signal analysis, enabling both common-noise rejection and field-normalized characterization of coherent response. The resulting spectra harbour the signatures of coherent interference between the resonant (vibrational) and nonresonant (electronic) contributions to the susceptibility, previously only observed under fs- and ps-pulsed excitation. We demonstrate the versatility and reproducibility of our approach across several molecules that yield distinct resonant and nonresonant mid-infrared responses. Together, these features position our platform as a scalable route toward multiplexed, high-resolution mid-IR sensing and a potential basis for coherent single- or few-molecule vibrational detection at the chip level.
Introduction
Spectroscopy in the mid-infrared (MIR) domain (400 to 4’000 cm-1) offers unique insights into molecular structures through their vibrational fingerprints [1]. Nonlinear optical techniques such as sum-frequency generation (SFG) and difference-frequency generation (DFG) combine a MIR pulse and a visible (VIS) pulse to probe vibrational modes that are both IR and Raman active, resulting in a resonant second-order () nonlinear response when the MIR light frequency matches a vibrational frequency. [2, 3, 4] Since bulk centro-symmetric media feature a vanishing coherent response, SFG and DFG are particularly useful to probe surfaces and interfacial vibrational modes [5, 6, 7]. They have been applied to study molecules bound to plasmonic nanoparticles [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29] and plasmonic antenna arrays [30, 31, 32, 33], revealing a wealth physico-chemical phenomena such as nanoparticle-induced alteration of molecular configurations [11, 13, 28], enhancement of the chiral molecule response [25], or plasmon-induced modification of vibrational relaxation [27, 29]. In all these experiments, the local field enhancement resulted in modest improvements in signal strength, so that the measurements had to be performed using femto- or picosecond-pulsed excitation schemes over macroscopic areas containing many plasmonic antennas or hot spots, imposing strict geometric phase-matching constraints. These requirements limit spectral tunability, spatial resolution, and sensor integration.
Here, using a versatile and broadband plasmonic nanocavity design, we demonstrate high-resolution ( cm-1, determined by the MIR laser tuning accuracy) and ultra-broadband (860 to 1670 cm-1) coherent MIR nonlinear spectroscopy using a commercial tunable continuous-wave (cw) MIR laser (MIRcat, Daylight Solutions) and a conventional low-resolution spectrometer (Shamrok 293, Andor) equipped with a silicon sensor. Our platform circumvents the need for geometric phase matching by confining light in the nanoparticle-on-slit (NPoS) construct that supports colocalized VIS and MIR resonances [34, 35]. The VIS resonances arise from the nanoparticle–substrate interaction and also support surface-enhanced Raman scattering (SERS) [36]. The MIR response is provided by the slit etched into the gold film, which acts as an inverted antenna and has previously been exploited for surface-enhanced infrared absorption (SEIRA) [37, 38, 39]. This dual-resonant configuration enables strong optical field confinement and enhancement within a deep-subwavelength volume, facilitating efficient nonlinear signal generation with relatively low-power continuous-wave excitation.
By arranging the infrared slit antennas into a metasurface, we can spectrally tune the MIR resonance to match different molecular vibrational modes. We illustrate this principle using two metasurface designs optimized for 1080 and 1600 cm-1, respectively, and test them with analytes such as biphenyl thiol (BPhT), 4-nitrothiophenol (4-NTP), and 4’-cyanobiphenyl-4-thiol (cn-BPhT). These molecules differ in their vibrational Raman and IR activities and electronic structures, offering a testbed for evaluating both spectral selectivity and signal reproducibility. The acquired spectra resolve multiple vibrational peaks from molecular monolayers embedded in individual NPoS cavities, yielding chemically specific signals with excellent signal-to-noise ratios across multiple structures and molecules.
The relaxed phase-matching condition at the single-cavity level allows for simultaneous acquisition of SFG and DFG signals, from which we compute the SFG/DFG ratio. This ratio-metric approach not only suppresses common-mode fluctuations, such as laser noise and sample drift, but also enables a field-normalized readout that is sensitive to molecular resonance features while suppressing baseline variations arising from the plasmonic resonances. Consistent SFG and DFG responses are observed across replicate structures and different analytes, supporting an interpretation based on coherent interference between resonant (vibrational) and non-resonant (electronic) contributions to the effective response.
In addition to its spectral range and resolution, the use of a cw tunable MIR laser enables measuring throughout the atmospheric water absorption window without purging (Fig. S3), thanks to the ability to avoid narrow water vapour lines during wavelength sweeps. The resulting spectra remain robust across a wide MIR range and require minimal environmental isolation. Our approach thus establishes a general and scalable platform for broadband, coherent nonlinear MIR spectroscopy under ambient conditions, compatible with existing commercial MIR microscopes such as those developed for optical photothermal infrared (O-PTIR) spectroscopy [40, 41]. This points toward future applications in label-free molecular fingerprinting [42], chip-integrated chemical analysis, and possibly few- or single-molecule coherent nonlinear spectroscopy [43], where ultrashort pulses often induce sample damage [44]. This work also lays the experimental groundwork for coherent, MIR-active platforms in nonlinear photonic metasurfaces [45, 46] and quantum molecular optomechanics [47].
Results
Principle of the experiment
The dual-resonant NPoS nanocavity [34] is an efficient geometry to boost SFG and DFG from a small number of molecules embedded in the nanogaps between the gold particle (150 nm nominal diameter) and the slit walls (Fig. 1a-c). The slit length determines the MIR resonance of each individual antenna [35], and the periodicity of a slit array can be leveraged to tune the overall MIR absorption spectrum (Fig. S4). Three examples are shown in Fig. 1d, overlaid with the Raman scattering spectrum of biphenyl-thiol (BPhT) collected from a single NPoS under 780 nm cw laser excitation. Each array is expected to offer broadband MIR field enhancement, with maxima around 1100 and 1600 cm-1, respectively. When the MIR laser is focused on the same NPoS through the silicon substrate by a reflective Cassegrain objective (NA 0.78, Pike Technologies), the second-order nonlinear response of the molecule and molecule-metal interface gives rise to simultaneous SFG and DFG, two coherent processes that are overlaid on the spontaneous incoherent anti-Stokes and Stokes Raman scattering signals, respectively (Figs. 1b,e). Two illustrative single-shot spectra (1 s exposure time) obtained under two MIR pump frequencies are presented in Fig. 1e. The net SFG and DFG signals are obtained by subtracting the spontaneous Raman signal (Fig. 1f) and are divided by the incident VIS and MIR powers for each MIR frequency (Fig. 1g).
Compared to SFG measurements performed on large-area plasmonic substrates under pulsed excitation, our method offers several advantageous features. First, the sub-wavelength dimension of the probed volume renders geometric phase-matching irrelevant, enabling simultaneous observation of SFG and DFG and straightforward tuning of the MIR frequency, limited only by the MIR laser range and the plasmonic resonance linewidth. We will show below that collecting both SFG and DFG helps constrain the phase of the resonant vs. non-resonant nonlinear responses, while their ratio is insensitive to common-mode noise, allowing for improved signal-to-noise. Second, operation under cw excitation provides very high spectral resolution (limited by the MIR laser tuning accuracy, which is better than 1 cm-1), irrespective of the employed spectrometer, offering opportunities for reducing the total system size and complexity.
Broadband, high-resolution SFG and DFG spectra from individual NPoS nanocavities
We discuss two sets of spectra obtained on individual nanocavities containing BPhT in Fig. 2, which are representative of many more measurements. Figs. 2a and b compare the data from nanocavities sitting on two slit arrays exhibiting distinct MIR absorption spectra. The first observation is that the overall strength of both SFG and DFG closely follows the MIR absorption spectrum of the slit array, confirming that both processes are mediated by the excitation of plasmonic resonances. More precisely, both absorption and frequency upconversion strengths are expected to scale with the square of the MIR field enhancement, explaining their similar spectra. Second, several sharp peaks and dips appear that are the expected signatures of vibrational resonances in the susceptibility interfering with the non-resonant response – a phenomenon that could not be revealed in previous experiments lacking broadband MIR coverage [34]. To better analyse these resonances, we plot in Fig. 2c the baseline-corrected SFG and DFG spectra as well as the SFG/DFG ratio. Remarkably, this ratio is self-corrected for the MIR absorption baseline since SFG and DFG share the same MIR frequency-dependent efficiency. Experimental data are shown as grey dots overlaid with analytical fit functions as full lines, obtained by expressing the nonlinear susceptibility as a coherent superposition of resonant and nonresonant components [2, 3, 4]:
| Symbol | Description |
| Amplitude of the -th vibrational mode response | |
| Phase offset of the -th mode | |
| Resonance frequency of the -th mode (fixed) | |
| Linewidth of the -th mode (fixed) | |
| Amplitude and phase of the nonresonant response | |
| Incident MIR frequency (experimentally controlled) |
The detected intensity is proportional to the squared modulus of the susceptibility:
We also evaluate the ratio:
In the non-resonant limit, the SFG/DFG ratio can be approximated as
so a flat baseline indicates nearly symmetric enhancement and similar non-resonant phases for SFG and DFG. Peaks or dips in (Fig. 2c) occur at frequencies where molecular vibrational modes generate a resonant response, with a relative phase offset that is mode-dependent.
Traditional spectral fitting of nonlinear optical signals such as SFG and DFG often relies on manual peak-wise adjustment or brute-force parameter sweeps. Such approaches become computationally infeasible or inaccurate in ultra-broadband regimes, where multiple overlapping resonances must be considered simultaneously. We employ a Bayesian Optimisation (BO) framework [48, 49] that jointly optimises the nonlinear parameters for SFG, DFG, and their ratio. This not only enhances fitting accuracy and speed but also enables global consistency across spectral channels, providing a more reliable physical model for ultra-broadband nonlinear spectroscopy. BO is a global optimisation method that is particularly effective for high-dimensional, expensive-to-evaluate problems. It builds a surrogate model (typically a Gaussian process) of the objective function and iteratively selects new parameter sets to evaluate by balancing exploration and exploitation, leading to faster and more accurate convergence.
For a vibrational mode to show up in SFG and DFG, it must be both IR and Raman active. To better evidence the differences between Raman and upconversion spectra, and to highlight the higher spectral resolution obtained by tuning the MIR laser compared to that provided by the VIS spectrometer, we show in Fig. 2d a zoomed-in view of the wavenumber region around 1’000 cm-1. We also present several independent measurements on different nanocavities to illustrate the robustness of the approach. What appears in Raman scattering as a peak doublet at 995 and 1’012 cm-1 hides a third vibrational mode close to 1’000 cm-1 with low Raman activity, which makes a large contribution to the SFG and DFG spectra due to its large IR activity. The existence of this strong IR vibrational peak is confirmed by density functional theory (DFT) calculations shown at the bottom of Fig. 2d. DFT, however, predicts a strong SFG mode at 1038 cm-1, which contrast with our experimental observation, showing the limits of molecular calculations in the context of complex nano-plasmonic environments. We tried to change the orientation of the molecule with respect to the gold plane in the simulations, but failed to identify a configuration that would reproduce the experimental data. Modifications of the vibrational lifetime due to the plasmonic near field have been evidenced in scanning-tip experiments [50] and may help explain the discrepancy with DFT; yet, it is unclear why other vibrational modes with strong IR dipoles would not be affected, such as the 1340 cm-1 mode of 4-NTP discussed below.
Applicability to other molecules
Finally, to validate the flexibility of our platform, we present Raman and broadband SFG spectra obtained with two other molecular species, namely, 4-nitrothiophenol (4-NTP, Figs. 3a,c) and 4’-cyanobiphenyl-4-thiol (cn-BPhT, Figs. 3b,d). Moreover, to bypass the restrictions associated with the formation of a uniform monolayer on the gold substrate, we instead functionalized the nanoparticles in solution before dropcasting them on the slit arrays. In the case of 4-NTP, the stability of the Raman signal is significantly less than for BPhT, obliging us to use lower VIS power of 1 W on the sample. Even in these challenging conditions, we obtain clean SFG and DFG signals: Figs. 3a,b show the raw spectra acquired on both molecules as a function of the MIR laser frequency, being swept over time from low to high wavenumbers. Compared to BPhT, the most striking feature, better visible in the net SFG spectra of Figs. 3c,d, is the much larger resonant to non-resonant contribution ratio for 4-NTP, resulting in strong simultaneous enhancement of SFG and DFG intensities when the MIR frequency matches the IR and Raman active vibration around 1’340 cm-1. This frequency has been shown to correspond to a collective vibrational mode mediated by large IR dipole-dipole coupling among neighbouring molecules in a self-assembled monolayer [51, 52]. cn-BPhT show an intermediate regime where resonant and non-resonant ’s are of the same order, leading to almost perfect cancellation of the SFG signal at some frequencies near 1’000 cm-1 (Fig. 3c). This happens when the vibrational has the same magnitude and opposite phase as the non-resonant electronic .
Discussion and perspectives
Several open questions remain. First, the SFG spectrum from BPhT presents significant deviations from that predicted by DFT calculations. Could this be related to the embedding of the molecules in a plasmonic nanogap? Is the discrepancy a signature of vacuum electromagnetic field modification [50], or rather linked to chemical interaction altering the geometry or electronic states of the molecules? Or is it related to other contributions not yet accounted for, such as from charge-transfer resonances? Unfortunately, we could not find broadband SFG spectra of BPhT from flat gold films in the literature to serve as a comparison. We attempted the DFT simulation of a more realistic geometry with molecules sandwiched between two gold planes using the model of Ref. [53], yet this did not yield better agreement with the data. Such complex metal-molecule interfaces are still a challenging system to accurately model with DFT.
Second, the strength of the non-resonant varies markedly compared to the resonant vibrational response, when changing the molecular species. This observation challenges the assumption that the non-resonant is an intrinsic property of the gold surface [54]. We speculate that the existence of charge-transfer resonances near the VIS laser wavelength at the gold-molecule interface may enhance the MIR non-resonant susceptibility, in the same way that it has been shown to contribute to Raman scattering enhancement in SERS [55, 56, 57]. Proving this hypothesis will require tuning the VIS wavelength over a broad range, as in [58, 54], or controllably shifting the energy of the charge-transfer state by molecular engineering [59] or application of electric fields [60] – which can also break inversion symmetry and thereby increase second-order nonlinearities [61]. An SFG study of dodecanethiol molecules adsorbed on Au, Ag, and Pt surfaces [58] suggested that the Fermi level of the substrate plays a key role in enhancing the non-resonant SFG signal. This enhancement could also be modulated by electron-withdrawing groups on the molecules.
All data presented above establish a reliable plasmonic platform combined with a simple laser technique to investigate the combined MIR-VIS nonlinear response of metallic nanogaps, which are taking central roles in a growing number of research and application fields, such as photo-catalysis [62], molecular and nano opto-electronics [63, 64], photodetection [65] and upconversion [66], nanoscale thermal transport [67, 68], polaritonics [69], vibrational strong coupling [70], 2D material opto-electronics [71], biosensors [72], chemical fingerprinting [73], single-molecule spectroscopy [74, 75], intra-molecular dynamics [76], etc. In Fig. S6, we show SFG spectra acquired in a single CCD exposure under fast continuous tuning of the MIR laser, a useful capability for applications requiring high throughput, such as hyperspectral far-field [40, 77, 78] or near-field [79, 80, 81] MIR and SFG imaging.
The platform’s versatility is validated across multiple molecules (BPhT, 4-NTP, cn-BPhT), showing reproducible, chemically specific signals from nanoscale volumes. The absence of geometric phase-matching requirements, the use of a tunable cw MIR laser, and the operation under ambient conditions pave the way for chip-integrated molecular sensing and single-molecule nonlinear spectroscopy. Given the high signal-to-noise ratio currently obtained with a few hundred molecules in the probed volume (see Fig. 1e), the detection of SFG from a single molecule seems within reach [82], using, for example, host-guest chemistry [83] or DNA origami techniques [84, 85] to trap a controlled number of molecules. In this regime, the anharmonicity of the molecular vibrations has been predicted to result in non-classical anti-bunching of upconverted photons [86], opening new applications in quantum optics and quantum-enhanced spectroscopy, not accessible to the recently developed single-molecule IR spectroscopy methods that leverage incoherent (e.g., fluorescence) readout [87, 88, 89, 90, 91, 92, 93, 94].
More generally, we foresee that our approach will enable new insights into intra-molecular [95, 96, 97] and inter-molecular [98, 99] vibrational relaxation pathways probed under plasmon-enhanced monochromatic MIR excitation. This includes the study of mode-specific, MIR-induced chemical reactions [100, 101], for which the SERS functionality built into our platform is a powerful way to track reactivity with single-molecule sensitivity [102]. The cavity-enhanced MIR excitation has also been predicted to enable Floquet engineering of molecular dynamics [103], resulting in a laser-tunable strength and linewidth of some vibrational sidebands.
Acknowledgments
F. C. thanks Roberto A. Boto for insightful exchanges about the DFT modelling. Z. X. and C. G. thank Matthias Godejohann for valuable assistance with the MIR laser, and the EPFL Cmi and IPHYS staff for technical support. This research received funding from the Swiss National Science Foundation (SNSF project No. 214993) and the European Union’s Horizon 2020 research and innovation program under Grant Agreement No. 820196 (ERC CoG ‘QTONE’)