### v2

# Continuous-wave, high-resolution, ultra-broadband mid-infrared nonlinear spectroscopy with tunable plasmonic nanocavities

Abstract
Vibrational sum- and difference-frequency generation (SFG and DFG) spectroscopy is a technique to probe the nonlinear response of interfaces at mid-infrared wavelengths while detecting the upconverted signal at visible wavelengths. It has recently transitioned from large-area molecular films and colloidal suspensions to nanoscale structures, enabled by dual-resonant plasmonic nanocavities that confine light and matter in the same deep-subwavelength sensing volume. Here, we implement high-resolution ( cm-1), continuous-wave ultra-broadband SFG and DFG spectroscopy from to on a variety of dual-resonant antennas under ambient conditions. The use of a commercial, broadly tunable quantum-cascade laser and the absence of geometric phase-matching requirements speed up and simplify data acquisition. At the same time, simultaneous access to SFG and DFG sidebands permits ratiometric signal analysis, enabling both common-noise rejection and field-normalized characterization of coherent response. The resulting spectra harbour the signatures of coherent interference between the resonant (vibrational) and nonresonant (electronic) contributions to the susceptibility, previously only observed under fs- and ps-pulsed excitation. We demonstrate the versatility and reproducibility of our approach across several molecules that yield distinct resonant and nonresonant mid-infrared responses. Moreover, we identify robust differences between our data and DFT calculations that hint at the modification of the infrared molecular response by the nanocavity environment. Together, these features position our platform as a scalable route toward multiplexed, high-resolution mid-IR sensing and a potential basis for coherent single- or few-molecule vibrational spectroscopy at the chip level.
Introduction
Coherent spectroscopy in the mid-infrared (MIR) domain (400 to 4’000 cm-1) offers unique insights into molecular structures through their vibrational fingerprints [1]. Nonlinear optical techniques, such as sum-frequency generation (SFG) and difference-frequency generation (DFG), combine an MIR pulse and a visible (VIS) pulse to probe vibrational modes that are both IR and Raman active. A resonant second-order () nonlinear response arises when the MIR frequency matches a molecular vibration [2, 3, 4]. In bulk centro-symmetric media, the coherent signal vanishes, making SFG and DFG especially valuable for probing surfaces and interfacial vibrational modes [5, 6, 7]. These techniques have been applied to study molecules bound to plasmonic nanoparticles [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29] and plasmonic antenna arrays [30, 31, 32, 33], revealing a wealth of physico-chemical phenomena, such as nanoparticle-induced alteration of molecular configurations [11, 13, 28], enhancement of the chiral molecule response [25], or modification of vibrational relaxation and dephasing dynamics [27, 34, 29, 35]. In most of these experiments, the local field enhancement provided only modest signal gains. Consequently, signal detection required femto- or picosecond-pulsed excitation schemes over macroscopic areas containing many plasmonic antennas or hot spots. This configuration, in turn, imposed strict geometric phase-matching conditions. These constraints limit spectral tunability, spatial resolution, and sensor integration.
Here, we demonstrate a coherent MIR nonlinear spectroscopy platform with sub-wavenumber ( cm-1) spectral resolution from 860 to 1670 cm-1, deep sub-wavelength spatial resolution, and few-molecule sensitivity. The approach uses a versatile plasmonic nanocavity design together with a commercial tunable continuous-wave (cw) MIR quantum cascade laser and a standard CCD spectrometer. Our platform circumvents the need for geometric phase matching by confining light in the nanoparticle-on-slit (NPoS) construct. This architecture simultaneously supports colocalized VIS and MIR resonances [36, 37], as also achieved using the nanoparticle-on-disk geometry [38, 35]. The VIS resonances arise from the nanoparticle–substrate interaction and enable surface-enhanced Raman scattering (SERS) [39]. The MIR response is provided by the slit etched into the gold film, which acts as an inverted antenna, and has previously been exploited for surface-enhanced infrared absorption (SEIRA) [40, 41, 42]. This dual-resonant configuration confines light to a deep-subwavelength volume, where strong optical field enhancement drives efficient nonlinear signal generation under relatively low-power cw excitation [36].
By adjusting the slits’ length and arranging them into a metasurface, arbitrary spectral tuning of the MIR resonance is achieved to match different molecular vibrational modes. We illustrate this principle using metasurfaces optimized for 1080 and 1600 cm-1, respectively, and test them with analytes such as biphenyl thiol (BPhT), 4-nitrothiophenol (4-NTP), and 4’-cyanobiphenyl-4-thiol (cn-BPhT). These molecules differ in their vibrational Raman and IR activities and electronic structures, offering a testbed for evaluating both spectral selectivity and signal reproducibility. The acquired spectra resolve multiple vibrational peaks from molecular monolayers embedded in individual NPoS cavities, yielding chemically specific signals with excellent signal-to-noise ratios across multiple structures and molecules.
Our platform natively captures SFG and DFG signals at the same time from the same cavity, whereas traditional methods must be complexified to achieve this [43]. This simultaneity ensures that both signals share identical MIR gain and noise. Calculating their ratio (SFG/DFG) suppresses common-mode fluctuations such as laser noise and sample drift. It also enables a field-normalized readout that isolates molecular resonance features while reducing baseline variations arising from the MIR plasmonic resonance. Consistent SFG and DFG responses are observed across replicate structures and different analytes, supporting an interpretation based on coherent interferences between resonant (vibrational) and non-resonant (electronic) contributions summing up to the effective response. Our experimental results also reveal clear and robust deviations from the DFT predictions of the SFG and DFG molecular signals, which motivate future work to clarify the impact of the nanocavity electromagnetic and chemical environment on the IR response of a molecular monolayer [44, 35].
In addition to its spectral range and resolution, the use of a cw tunable MIR laser allows measuring across the atmospheric water absorption window without purging, by simply avoiding narrow water vapour lines during wavelength sweeps. This capability yields robust spectra under ambient conditions and requires minimal environmental isolation. Together, these features establish a general and scalable platform for broadband coherent nonlinear MIR spectroscopy, directly compatible with commercial MIR microscopes such as those developed for O-PTIR [45, 46] (Optical photothermal infrared spectroscopy). The approach presented here opens opportunities in label-free molecular fingerprinting [47], chip-integrated chemical analysis, and few- or single-molecule coherent nonlinear spectroscopy [48], where ultrashort pulses often induce sample damage [49]. More broadly, it extends the domain of nonlinear plasmonics beyond the UV-VIS and near-infrared wavelengths [50, 51, 52, 53, 54, 55] to provide an experimental foundation for low-power MIR-active nonlinear metasurfaces [56, 57], quantum molecular optomechanics [58], and the study of inter- and intra-molecular vibrational energy transfer in the vibrational strong coupling regime [59, 60].
Results
Principle of the experiment
The dual-resonant NPoS nanocavity [36, 37] is an efficient geometry to boost SFG and DFG from a small number of molecules embedded in the nanogaps between the gold particle (150 nm nominal diameter) and the slit walls (Fig. 1a-b). We fabricate densely packed arrays of nanoslits acting as MIR metasurfaces according to the process described in Fig. S1. The slit length determines the MIR resonance of individual antennas [37], and the periodicity of a slit array can be leveraged to tune the overall MIR absorption spectrum, as measured by conventional Fourier-transform IR spectroscopy (FTIR, Fig. S2). Three examples of MIR metasurface absorption spectra are shown in Fig. 1c, overlaid with the Raman scattering (SERS) spectrum of biphenyl-thiol (BPhT) collected from a single NPoS under 780 nm cw laser excitation. Importantly, the SERS spectrum is found to be nearly identical on all NPoS probed in our study, irrespective of the metasurface parameters. When the MIR laser is focused on the same NPoS through the silicon substrate by a reflective Cassegrain objective (NA 0.78, Pike Technologies; see Fig. S3), the molecule and the molecule-metal interface produce a second-order nonlinear response. This response gives rise to simultaneous SFG and DFG signals, which are overlaid on the spontaneous anti-Stokes and Stokes Raman scattering signals, respectively (Figs. 1b,d).
Two illustrative single-shot (1 s exposure time) emission spectra (Raman+SFG+DFG) obtained under two MIR pump frequencies and the same VIS pump condition are presented in Fig. 1d. In this plot, we use the Raman spectroscopy convention, whereby a negative Raman shift corresponds to anti-Stokes (higher frequency) emission. The net SFG and DFG signals are obtained by subtracting the spontaneous Raman signal from the total emission (Fig. 1e) and are divided by the product of incident VIS and MIR powers, for each MIR frequency (see Fig. S4 for the measured calibration curve). A subset of the resulting normalized DFG peaks is presented in Fig. 1f. The MIR frequency can be sweept with increments down to 0.5 cm-1 (specified tuning accuracy of the MIRcat laser), and we often use 1 cm-1 steps. In the following, we will display SFG and DFG spectra that represent the normalized SFG and DFG peak intensity vs. MIR frequency (expressed in wavenumber).
Broadband, high-resolution SFG and DFG spectra from individual NPoS nanocavities
We discuss two sets of SFG and DFG spectra obtained on individual nanocavities containing BPhT in Fig. 2, which are representative of many more measurements (see Figs. S5, S6). Figures 2a and b compare the data from typical NPoS on two metasurfaces (M-1 and M-2) exhibiting distinct MIR absorption spectra. The first observation is that the overall SFG and DFG intensities track the MIR absorption spectrum of the slit array, confirming that the excitation of plasmonic resonances mediates both processes [37]. More precisely, both absorption and frequency upconversion strengths are expected to scale with the square of the MIR field enhancement, which explains their similar spectral profiles. Second, several sharp peaks and dips emerge. This is consistent with the expected signatures of vibrational resonances in the susceptibility, which interfere with the non-resonant background. This phenomenon could not be observed in previous experiments on similar structures because they lacked broadband MIR coverage [36] or exhibited incoherent response or insufficient resolution [38].
To better analyse these resonances, we plot in Fig. 3a the baseline-corrected SFG and DFG spectra as well as the self-corrected SFG/DFG ratio. Experimental data are shown as grey dots overlaid with analytical fit functions as full lines, obtained by expressing the nonlinear susceptibility as a coherent superposition of resonant and nonresonant components, where vibrational frequencies and linewidths are fixed by the experimental Raman spectra, as noted in Table 1 [2, 3, 4]:
| Symbol | Description |
| Amplitude of the -th vibrational mode response | |
| Phase offset of the -th mode | |
| Resonance frequency of the -th mode (fixed) | |
| Linewidth of the -th mode (fixed) | |
| Amplitude and phase of the nonresonant response | |
| Incident MIR frequency (experimentally controlled) |
A complete expression of the nonlinear signal strength is discussed in Ref. [37]. It involves the vectorial field overlap among all involved frequencies, mediated by the -tensor elements. Here, eqs. (1,2) describe only the frequency dependence of the effective (contracted) around each vibrational resonance. As the MIR and VIS plasmonic resonances are much broader, the mode overlap is expected to vary slowly with frequency, contributing to the baseline shape. Under these assumptions, the detected intensity is proportional to the squared modulus of the effective nonlinear susceptibility:
We also evaluate the ratio:
Away from vibrational resonances, the ratio reduces to a constant. Consequently, a flat baseline indicates a nearly constant mode overlap across the probed MIR range. Peaks or dips in (Fig. 3c) occur at frequencies where molecular vibrational modes generate a resonant response, with a relative phase offset that is mode-dependent.
Traditional fitting of nonlinear spectra often relies on manual peak adjustment or brute-force parameter sweeps, which become computationally infeasible or inaccurate in ultra-broadband regimes with multiple overlapping resonances. We instead use a Bayesian Optimisation (BO) framework [61, 62] that jointly optimises the nonlinear parameters for SFG, DFG, and their ratio. This apprroch accelerates convergence, improve fitting accuracy and enables global consistency across spectral channels, providing a more reliable physical model for ultra-broadband nonlinear spectroscopy. BO builds a surrogate model (typically a Gaussian process) of the objective function, and iteratively selects parameter sets that balance exploration and exploitation, leading to faster and more accurate convergence.
For a vibrational mode to show up in SFG and DFG, it must be both IR and Raman active. To better evidence the differences between Raman and upconversion spectra, and to highlight the higher spectral resolution obtained by tuning the MIR laser compared to that provided by the VIS spectrometer, we show in Fig. 3b a zoomed-in view of the SFG/DFG ratio over the wavenumber region from 970 to 1’110 cm-1. We also present several independent measurements on different NPoS from M1 and M2 metasurfaces to illustrate the robustness of the approach. In Raman scattering, the spectral feature close to 1’000 cm-1 apears as a doublet at 995 and 1’012 cm-1. However, our SFG and DFG reveal a third vibrational mode at 1’003 cm-1 with weak Raman but strong IR activity, in agreement with density functional theory (DFT) calculations shown at the bottom of Fig. 3b (see section S6 for details about the DFT model).
DFT also predicts two strong SFG-active modes at 1038 cm-1 and 1’080 cm-1, which appear much weaker experimentally. Adjusting the orientation of the molecule with respect to the gold plane in the simulations did not reconcile this discrepancy, highlighting the limits of molecular calculations in complex plasmonic enviroments. One possible explanation is a modification of the vibrational lifetime due to the enhanced MIR plasmonic near field, as previously proposed in scanning-tip experiments [44]. The Purcell effect, if strong enough, can indeed cause a lifetime broadening of the IR-active transition, following the same physics as optically-allowed electronic transition of molecules at VIS frequencies [63]. However, it remains unclear why other vibrational modes with strong IR dipoles, such as the 1340 cm-1 mode of 4-NTP discussed below, are not similarly affected. These discrepancies underscore the difficulties of state-of-the-art DFT models to capture complex metal–molecule and plasmon-vibration interactions. Nevertheless, our experimental SFG/DFG spectra are robust across molecules and metasurfaces, preserving line shapes while suppressing MIR baseline drifts. Combined with the compact cw implementation, this capability positions the platform as a practical bridge between chip-integrated photonics and chemically specific MIR spectroscopy that is sensitive to local chemical and photonic environments.
Figs. 2 and 3 show that our method offers several advantageous features compared to pulsed SFG measurements performed on large-area plasmonic substrates: (i) The sub-wavelength probing volume renders geometric phase-matching irrelevant, enabling simultaneous observation of SFG and DFG and straightforward tuning of the MIR frequency, limited only by the MIR laser range and the plasmonic resonance linewidth. (ii) Collecting both SFG and DFG helps extract the phase of the resonant vs. non-resonant nonlinear responses, while their ratio is insensitive to common-mode noise, allowing for improved signal-to-noise. (iii) Operation under cw excitation provides very high spectral resolution (limited by the MIR laser tuning accuracy, which is better than 1 cm-1), irrespective of the employed spectrometer, offering opportunities for reducing the total system size and complexity. (iv) By probing a deep-subwavelength area of the molecular layer, we unveil spatial variations in the coherent signal that appear to be more sensitive to the local environment than the SERS signal.
Applicability to other molecules
Finally, to validate the flexibility of our platform, we present SERS and broadband SFG spectra obtained with two other molecular species, namely, 4’-cyanobiphenyl-4-thiol (cn-BPhT, Figs. 4a), and 4-nitrothiophenol (4-NTP, Figs. 4b). The SFG and DFG spectra from cn-BPhT are even more clearly marked by the constructive and destructive interferences between resonant (vibrational) and non-resonant (electronic) responses. This means that these two terms have similar magnitudes in this case, whereas the non-resonant response dominated for BPhT.
In the case of 4-NTP, the stability of the Raman signal is significantly less than for BPhT because of its photochemical reactivity [64, 65, 66, 67, 68], obliging us to use lower the VIS power down to 1 W on the sample. Even in these challenging conditions, we obtain clean SFG and relatively weaker DFG signals (potentially due to a change in the VIS plasmonic resonances compared to BPhT NPoS): Fig. 4b show the raw spectra acquired on 4-NTP molecules as a function of the MIR laser frequency, being swept over time from low to high wavenumbers. Compared to BPhT, 4-NTP shows a much larger resonant to non-resonant ratio when the MIR frequency matches the IR and Raman active vibration around 1’340 cm-1, as seen in the net SFG spectra of Figs. 4b. Previous studies have attributed this frequency to a collective vibrational mode mediated by strong IR dipole-dipole coupling between neighbouring molecules within a self-assembled monolayer [69, 70, 35].
The comparisons across BPhT, 4-NTP, and cn-BPhT emphasize that our platform can capture different nonlinear regimes: With BPhT, the non-resonant dominates the SFG and DFG response over most of the measurement bandwidth, while the signal from 4-NTP mostly comes from the vibrational resonances, which appear as symmetric peaks in SFG and DFG. An intermediate regime is observed for cn-BPhT. This capability is essential for disentangling spatially-dependent molecular contributions to that are obscured in ensemble measurements and will be useful for investigating the impact of nanoscale symmetry breaking on nonlinear conversion efficiency [55].
To position our cw-SFG/DFG nanocavity-enhanced platform within the broader MIR spectroscopy landscape, we benchmark it against representative approaches using literature values (Table 2). This comparison underscores in which scenarios our platform may outperform other methods, namely, when a combination of few-molecule sensitivity and a simplified, compact apparatus is needed, while maintaining the highest possible spectral resolution and molecular specificity through coherent spectroscopy. Ultimately, the quantum cascade laser will drive the size and complexity of the system, since there is no need for a spectrometer in the detection path. To support this claim, we emulate in Fig. 4b (middle panel) the SFG spectrum that would be obtained using a single-pixel detector behind a bandpass filter that selects the window of interest on the anti-Stokes side. In this hypothetical scenario, the readout of a sensitive detector, such as an avalanche photodiode, is demodulated at the QCL modulation frequency (on the order of a MHz) to recover the differential (‘MIR ON’ minus ‘MIR OFF’) spectrally integrated signal. We simulate the result by summing the intensity of all CCD pixels within the spectral range from 870 to 1650 cm-1, and obtain a spectrum that is almost identical to that of the top panel, derived from the refined analysis presented before in Fig. 1.
| Method | Coherent | Simult. SFG+DFG | Standalone MIR source | Label-freea | Number of molecules |
| This work | Yes | Yes | Yes | Yes | |
| Nanocavity pulsed SFG [35] | Yes | Possible | No | Yes | |
| Free-space pulsed SFG [3] | Yes | Possible but complex | No | Yes | typ. |
| O-PTIR [71, 46] | No | (NA) | Yes | Yes | |
| BonFIRE (VEF) [72] | No | (NA) | No | No | single |
| MIRVAL (VEF) [73] | No | (NA) | No | No | c |
| SEIRA on a SAM [74] | No | (NA) | Yes | Yes | typ. |
| nano-FTIRb [75] | No | (NA) | Yes | Yes |
Discussion
Several open questions remain. First, the SFG spectrum from BPhT presents significant deviations from that predicted by DFT calculations. Could this be related to the embedding of the molecules in a plasmonic nanogap? Is the discrepancy a signature of vacuum electromagnetic field modification [44], or rather linked to chemical interaction altering the geometry or electronic states of the molecules? Or is it related to other contributions not yet accounted for, such as from charge-transfer resonances? Unfortunately, we could not find broadband SFG spectra of BPhT from flat gold films in the literature to serve as a comparison. We attempted the DFT simulation of a more realistic geometry with molecules sandwiched between two gold planes using the model of Ref. [76], yet this did not yield better agreement with the data. Such complex metal-molecule interfaces are still challenging systems to accurately model with DFT.
Second, the strength of the non-resonant varies markedly compared to the resonant vibrational response when changing the molecular species. This observation challenges the assumption that the non-resonant is an intrinsic property of the gold surface [77]. We speculate that the existence of charge-transfer resonances near the VIS laser wavelength at the gold-molecule interface may enhance the MIR non-resonant susceptibility, in the same way that it has been shown to contribute to Raman scattering enhancement in SERS [78, 79, 80]. Proving this hypothesis will require tuning the VIS wavelength over a broad range, as in [81, 77], or controllably shifting the energy of the charge-transfer state by molecular engineering [82] or application of electric fields [83] – which can also break inversion symmetry and thereby increase second-order nonlinearities [84]. An SFG study of dodecanethiol molecules adsorbed on Au, Ag, and Pt surfaces [81] suggested that the Fermi level of the substrate plays a key role in enhancing the non-resonant SFG signal. This enhancement could also be modulated by electron-withdrawing groups on the molecules.
Conclusion and perspectives
All data presented above establish a reliable plasmonic platform combined with a simple laser technique to investigate the combined MIR-VIS nonlinear response of metallic nanogaps. Such platforms are taking central roles in a growing number of research and application fields, such as photo-catalysis [85], molecular and nano opto-electronics [86, 87], direct photodetection [88], frequency upconversion [89, 38, 90, 91], nanoscale thermal transport [92, 93], polaritonics [94], vibrational strong coupling [95], 2D material opto-electronics [96], biosensors [97], chemical fingerprinting [98], single-molecule spectroscopy [99, 100], intra-molecular dynamics [101], etc. In Fig. S6, we show SFG spectra acquired in a single CCD exposure under fast continuous tuning of the MIR laser, a useful capability for applications requiring high throughput, such as hyperspectral far-field [45, 102, 103] or near-field [104, 105, 106] MIR and SFG imaging.
The platform’s versatility is validated across multiple molecules (BPhT, 4-NTP, cn-BPhT), showing reproducible and chemically specific signals from nanoscale volumes. Importantly, the absence of geometric phase-matching requirements, the use of a tunable cw MIR laser, and the operation under ambient conditions pave the way for chip-integrated molecular sensing.Given the high signal-to-noise ratio currently obtained with a few hundred molecules in the probed volume (see Fig. 1e), the detection of SFG from a single molecule seems within reach [107] (see Fig. S7). This could be realized using host-guest chemistry [108] or DNA origami techniques [109, 63] to trap a controlled number of molecules. In this regime, the anharmonicity of the molecular vibrations has been predicted to result in non-classical anti-bunching of upconverted photons [110]. Such quantum signatures would open new applications in quantum optics and quantum-enhanced spectroscopy, which are not accessible to recently developed single-molecule IR spectroscopy methods based on incoherent (e.g., fluorescence) readout [111, 112, 71, 113, 73, 72, 114, 115, 116].
More generally, we foresee that our approach will enable new insights into intra-molecular [117, 118, 60] and inter-molecular [59, 119] vibrational relaxation pathways probed under plasmon-enhanced monochromatic MIR excitation. This includes the study of mode-specific, MIR-induced chemical reactions [120, 121], for which the SERS functionality built into our platform is a powerful way to track reactivity with single-molecule sensitivity [122]. The cavity-enhanced MIR excitation has also been predicted to enable Floquet engineering of molecular dynamics [123], resulting in a laser-tunable strength and linewidth of some vibrational sidebands.
Acknowledgments
F. C. thanks Roberto A. Boto for insightful exchanges about the DFT modelling. Z. X. and C. G. thank Matthias Godejohann for valuable assistance with the MIR laser, the EPFL Cmi and IPHYS staff for technical support, and Arnaud Jollivet for earlier contributions to the setup. This research received funding from the Swiss National Science Foundation (SNSF project No. 214993) and the European Union’s Horizon 2020 research and innovation program under Grant Agreement No. 820196 (ERC CoG ‘QTONE’)